OpenStudy (gabylovesyou):

Andrew opened a car wash station in his neighborhood. He recorded the number of cars washed each week for the first six weeks. The data are shown as a scatter plot below.

OpenStudy (gabylovesyou):

Based on the line of best fit, which of these would be the best prediction for the number of cars washed in Week 8? 4 9 16 18

OpenStudy (gabylovesyou):

@SmoothMath

OpenStudy (anonymous):

can you find the equation of the line ?

OpenStudy (gabylovesyou):

how?

OpenStudy (anonymous):

you have a point 0,0 on the graph and 4,9

OpenStudy (anonymous):

Not even necessary. Just look at where the line is going. What y value does it make sense that it would be at for week 8?

OpenStudy (anonymous):

using two points .. you find the slope m = (y1 - y2) /(x1-x2) = (9-0)/(4-0) = 2.25 now in order to get the equation of the line : y = y1 + m(x-x1) we choose one of the points so we choose 0,0 since its easier to work with y = 0 + 2.25(x-0) y = 2.25x in week 8 there will be y = 2.25 * 8 = 18